A bricklayer and an electrician together spend 90 hours working on a new house. If the bricklayer earns $12 per hour, the electrician earns $16 an hour, and the owner pays them a total of $1350 for their work, how many hours does each worker spend on the house?
Answer by checkley71(8403)

x+y=90 or x=90-y
12x+16(90-x)=1350
12x+1440-16x=1350
-4x=1350-1440
-4x=-90
x=-90/-4
x=22.5 hours for the bricklayer.
90-22.5=67.5 hours for the electrician.
proof
12*22.5+16*67.5=1350
270+1080=1350
1350=1350